Saint Joseph Travel and Tours Inc. borrows ₱ 45,000 from the bank for one year at an annual rate of 8% compounded quarterly. 1. How much interest will the company pay on this loan? 2. What is the effective rate of the loan?
1
Expert's answer
2021-10-27T19:04:07-0400

1.

"FV=PV(1+r)^n=45000(1+0.08\/4)^{4}=48709.45"

company will pay interest:

"48709.45-45000=3709.45"


2.

the effective rate of the loan:

"(1+0.08/4)^4-1=0.0824=8.24\%"